Solve the following equations. 
17 + 6p = 9 
Solve the following equations. 
17 + 6p = 9 

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

17 + 6p =9
6p = 9–17
6p = -8
p = -8/6
p = -4/3

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve it as follows ,
17+6p = 9
6p = 9 - 17
6p = -8
p = -8/6
P = -4/3
 
Hope it helps!
Thank you ✌️